OpenStudy (anonymous):

Check my work please... \[\frac{ -18a ^{-8}b ^{-3} }{ 6a ^{-2}b ^{-7} }\]

OpenStudy (anonymous):

-18 / 6 is -3 \[a ^{\left( -8 - (2) \right)} = a ^{-6}\]

OpenStudy (anonymous):

and \[b ^{\left( -3 - (-7) \right)} = b ^{4}\]

OpenStudy (anonymous):

So then \[\frac{ -3b ^{4} }{ a ^{6} }\]?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thanks!